. The equation which cannot be solved in integers is (a) 5y – 3 = – 18 (b) 3x – 9 = 0 (c) 3z + 8 = 3 + z (d) 9y + 8 = 4y – 7

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Option (c) is correct here

Consider the equation 3z + 8 = 3 + z
Transpose 8 to right hand side then it becomes–8 and z to Left hand side then it becomes - z
3z–z = 3–8
2z = -5
z = -5/2

Thanks